# BEVTrack: A Simple and Strong Baseline for 3D Single Object Tracking in Bird’s-Eye View

### v5

@@ I Introduction
3D single object tracking (SOT) is crucial for various applications, such as autonomous driving [1, 2, 3]. It aims to localize a specific target across a sequence of point clouds, given only its initial status. Existing tracking approaches [4, 5] commonly adopt the point-wise representation, directly taking raw point clouds as input. For example, P2B[4] and its follow-up works [6, 7, 8] adopt a point-based network [9, 10] with the Siamese architecture for feature extraction, followed by a point-wise appearance matching module [4, 8, 11] for propagation of target cues, and a 3D Region Proposal Network [12, 13] for target localization, as illustrated in Fig. 1(a). M2-Track[5] proposes a motion-centric paradigm, that first segments the target points from their surroundings with a PointNet [14] segmentation network and then localizes the target through a motion modeling approach followed by a box refinement module, as illustrated in Fig. 1(b).
@@
Although these approaches have exhibited superior performance in tracking benchmarks, two deficiencies still exist. First, they usually require elaborate designs [11, 15, 8, 16], leading to a complicated tracking framework. Second, they usually require solving multiple subtasks in the training phase, which makes the training cumbersome due to heavy hyperparameter tuning.
In this paper, we present BEVTrack, a simple yet strong baseline for 3D SOT, as shown in Fig. 1(c). By estimating the target motion in Bird’s-Eye View (BEV) to perform tracking, BEVTrack demonstrates surprising simplicity from various aspects, i.e., network designs, training objectives, and tracking pipeline, while achieving superior performance. It is based upon the intuition that BEV could better capture motion features, since objects primarily move horizontally in autonomous driving scenarios. Specifically, we first adopt a voxel-based network [17] with the Siamese architecture for feature extraction. Subsequently, we compress height information into the channel dimension to obtain the BEV features. Given that corresponding objects are spatially adjacent in the BEV features across consecutive frames, we can easily fuse them together with an element-wise operation such as concatenation [18, 19]. Then we adopt several convolutional and downsampling layers to capture the object motion with a wide range of patterns. Finally, we discard the complicated region proposal network while using a global max pooling followed by a lightweight multilayer perceptron (MLP) to localize the target, resulting in a simple one-stage tracking pipeline with a single regression loss.
@@ II Related Work
Early 3D SOT approaches [23, 24] predominantly utilize RGB-D data and often adapt 2D Siamese networks [25, 26, 27, 28] by integrating depth maps. However, variations in illumination and appearance can adversely affect the performance of such RGB-D techniques. SC3D [29] is the first 3D Siamese tracker based on shape completion that generates a large number of candidates in the search area and compares them with the cropped template, taking the most similar candidate as the tracking result. The pipeline relies on heuristic sampling and does not learn end-to-end, which is very time-consuming.
@@
Some prior research also utilizes BEV representations for tracking. Zarzar et al. [36] leverage the dense and structured BEV representation of LIDAR point clouds to efficiently search for objects of interest. LTTR [37] locates targets through similarity matching of BEV features. V2B [38] recommends converting the enhanced point features into dense BEV feature maps to generate high-quality target proposals. Recently, STTracker [39] fuses multi-frame BEV features to enhance the temporal features. These methods are based on appearance matching, with an RPN network [12, 40] for localization. In contrast, our BEVTrack is matching-free, and directly regresses the relative target motion in BEV for the first time.
@@
3D object detection methods are usually similar to their 2D counterparts, such as R-CNN series [41, 42, 43, 44, 45] and CenterPoint series [46, 40, 47]. VoxelNet [17] utilizes PointNet [14] for voxel feature encoding, and then converts 3D sparse features into 2D dense BEV features [48]. Finally, it adopts a region proposal network for prediction. SECOND [19] improves VoxelNet by proposing efficient sparse convolutions. Motivated by 2D CenterNet [46], CenterPoint [40] focuses on 3D detection and tracking. It also transforms the sparse output from a backbone network into a dense feature map, enabling the prediction of dense heatmaps indicating the center locations of objects. Recently, VoxelNext [49] predicts objects directly based on sparse voxel features, realizing fully sparse 3D object detection. In this paper, we adopt a variant of VoxelNet [17] as the backbone, which modifies the original architecture by removing residual connections, utilizing sparse 3D convolutions, and choosing suitable network hyper-parameters, such as voxel size.
@@
3D multi-object tracking models tracklets of multiple objects along multi-frame LiDAR. The majority of prior works in 3D MOT leverage the LiDAR modality. Due to the recent advances in LiDAR-based 3D detection [50, 49, 45], especially the reliable range information, most state-of-the-art 3D MOT algorithms [51, 52] adopt a “tracking-by-detection” paradigm. After detecting objects in each frame, these methods perform heuristic association of detected bounding boxes based on object motion or appearance [53, 54, 55, 56, 57]. CenterPoint [40] predicts the velocities to associate object centers through multiple frames, following CenterTrack [53]. SimTrack [58] combines object detection and motion association in a unified pipeline for joint detection and tracking. Compared to the 3D MOT methods, our BEVTrack focuses on tracking a specific target of interest.
@@ V Conclution
Limitation: One limitation of BEVTrack is that, despite being robust to typical challenging scenarios, it tends to drift to background points when point clouds are extremely sparse. A promising solution is to incorporate multi-modal information. On the one hand, image data can complement the sparsity of point cloud data. On the other hand, BEV, as the unified representation for multi-modal fusion [65, 66, 67], adeptly preserves semantic information from cameras and geometric information from LiDAR. We envision that BEVTrack can serve as an inspiration for further research into multi-modal fusion methodologies for SOT.

### v6

@@ I Introduction
3D single object tracking (SOT) is crucial for various applications, such as autonomous driving [1, 2]. It aims to localize a specific target across a sequence of point clouds, given only its initial status. Existing tracking approaches [3, 4] commonly adopt the point-wise representation, directly taking raw point clouds as input. For example, P2B[3] and its follow-up works [5, 6, 7] adopt a point-based network [8, 9] with the Siamese architecture for feature extraction, followed by a point-wise appearance matching module [3, 7, 10] for propagation of target cues, and a 3D Region Proposal Network [11, 12] for target localization, as illustrated in Fig. 1(a). M2-Track[4] proposes a motion-centric paradigm, that first segments the target points from their surroundings with a PointNet [13] segmentation network and then localizes the target through a motion modeling approach followed by a box refinement module, as illustrated in Fig. 1(b).
@@
Although these approaches have exhibited superior performance in tracking benchmarks, two deficiencies still exist. First, they usually require elaborate designs, leading to a complicated tracking framework. Second, they usually require solving multiple subtasks in the training phase, which makes the training cumbersome due to heavy hyperparameter tuning.
In this paper, we present BEVTrack, a simple yet strong baseline for 3D SOT, as shown in Fig. 1(c). By estimating the target motion in Bird’s-Eye View (BEV) to perform tracking, BEVTrack demonstrates surprising simplicity from various aspects, i.e., network designs, training objectives, and tracking pipeline, while achieving superior performance. It is based upon the intuition that BEV could better capture motion features, since objects primarily move horizontally in autonomous driving scenarios. Specifically, we first adopt a voxel-based network [14] with the Siamese architecture for feature extraction. Subsequently, we compress height information into the channel dimension to obtain the BEV features. Given that corresponding objects are spatially adjacent in the BEV features across consecutive frames, we can easily fuse them together with an element-wise operation such as concatenation [15]. Then we adopt several convolutional and downsampling layers to capture the object motion with a wide range of patterns. Finally, we discard the complicated region proposal network while using a global max pooling followed by a lightweight multilayer perceptron (MLP) to localize the target, resulting in a simple one-stage tracking pipeline with a single regression loss.
@@ II Related Work
Early 3D SOT approaches [19, 20] predominantly utilize RGB-D data and often adapt 2D Siamese networks by integrating depth maps. However, variations in illumination and appearance can adversely affect the performance of such RGB-D techniques. SC3D [21] is the first 3D Siamese tracker based on shape completion that generates a large number of candidates in the search area and compares them with the cropped template, taking the most similar candidate as the tracking result. The pipeline relies on heuristic sampling and does not learn end-to-end, which is very time-consuming.
@@
3D object detection methods are usually similar to their 2D counterparts, such as R-CNN series [29, 30, 31, 32, 33] and CenterPoint series [34, 1]. VoxelNet [14] utilizes PointNet [13] for voxel feature encoding, and then converts 3D sparse features into 2D dense BEV features [35]. Finally, it adopts a region proposal network for prediction. Motivated by 2D CenterNet [34], CenterPoint [1] focuses on 3D detection and tracking. It also transforms the sparse output from a backbone network into a dense feature map, enabling the prediction of dense heatmaps indicating the center locations of objects. Recently, VoxelNext [36] predicts objects directly based on sparse voxel features, realizing fully sparse 3D object detection. In this paper, we adopt a variant of VoxelNet [14] as the backbone, which modifies the original architecture by removing residual connections, utilizing sparse 3D convolutions, and choosing suitable network hyper-parameters, such as voxel size.
@@
3D multi-object tracking models tracklets of multiple objects along multi-frame LiDAR. The majority of prior works in 3D MOT leverage the LiDAR modality. Due to the recent advances in LiDAR-based 3D detection [36, 37, 38], especially the reliable range information, most state-of-the-art 3D MOT algorithms [39, 40] adopt a “tracking-by-detection” paradigm. After detecting objects in each frame, these methods perform heuristic association of detected bounding boxes based on object motion or appearance [41, 42, 43, 44]. CenterPoint [1] predicts the velocities to associate object centers through multiple frames, following CenterTrack [41]. SimTrack [45] combines object detection and motion association in a unified pipeline for joint detection and tracking. Compared to the 3D MOT methods, our BEVTrack focuses on tracking a specific target of interest.
@@ IV-C Ablation Studies
Figure 8: Tracking failure cases of our BEVTrack on extremely sparse point cloud scene.
IV-D Limitation
One limitation of BEVTrack is that, despite being robust to typical challenging scenarios, it tends to drift to background points when point clouds are extremely sparse. As shown in Fig. 8, though the target reappears in subsequent frames, the model has mistaken background points as the target, resulting in failures. A promising solution is to incorporate multi-modal information. On the one hand, image data can complement the sparsity of point cloud data. On the other hand, BEV, as the unified representation for multi-modal fusion [53, 54], adeptly preserves semantic information from cameras and geometric information from LiDAR. We envision that BEVTrack can serve as an inspiration for further research into multi-modal fusion methodologies for SOT.